QUESTION 12 points Save Ans Calculate the force in member DJ of the Howe roof truss ilustrated below 10 kN D 10 kN E 6 m F K H Y10 KN -6 panels at 4 m- 10.17 kN,T 11.89 KN, C 637 kN, C 16.67 kN, T
QUESTION 13 5 points For the truss shown in Fig. below if the lifting force of 120 kN moves at a constant velocity of 8 m per sec. 8m 8m 120 kN The force in member (BC) is: 125 kN 204 kN O 192 kN
Answer #1

12)sing merthod f Sec tiens 0KN DR /oKN 10x 20 10x16iox 8- RIXL4-0 R18.33N R2t 13. 33- 10 l0-lo0 R2 167 KN FBo fsecfen (om CK2MA 0 CJ Sin45 -lox4-10x8 0 CT14.14KN (C) B= tari 26.5 12 lox4+lox8- CD COS26x6-13.33 X12D -366 C0 99.96 .CD= 13.63KNCC) FBD4 P=120 KN 120KN 120t120(3) C192KN 2FH-O CH=120(4)=96EN Gm By Symmeky of veial force A 966 8m RA Dy RA D96KN EFH=0 = Dこ96 ky